OpenStudy (junyang96):

find the strain Ez of a cube under constant volume deformation. Given strain Ex=30%,Ey=35%

OpenStudy (irishboy123):

you should prolly post your source formula but from \(V = xyz\) you have via under "constant volume deformation": \(\delta V = \color{red}{0} = \delta x . yz + \delta y .xz + \delta z. xy \qquad \large \color{red}{\star} \) and then: \(\dfrac{\delta x}{x} = 0.3\) and \(\dfrac{\delta y}{y} = 0.35\) then \(\) so, from \(\color{red}{\star}\) you have \(\dfrac{\delta z}{ z} \left[ =E_z\right] = -\dfrac{\delta x}{x} - \dfrac{\delta y}{y}\)

OpenStudy (irishboy123):

though that is highly simplistic under the conditions you specify

OpenStudy (junyang96):

That's what I thought. -65% but the answer given to me is 35% . Here's the original question

OpenStudy (anonymous):

k